The royal fruit company produces two types of fruit drinks. The first type is 45% pure fruit juice, the second type is 100% pure fruit juice. The company is attempting to produce a fruit drink that contains 80% pure fruit juice. How many pints of each of the existing types of drink must be used to make 110 pints of mixture that is 80% pure fruit juice?
Answer by addingup(3677)

Total 110 pints at a concentration of 0.80
first type at 0.45: x pints
second at 1 : 110-x
:
0.45x+1(110-x) = 0.8(110)
0.45x+110-1x = 88
-0.55x = -22 divide both sides by 0.55 and remember -/- = + so when you divide on your calculator just divide like positive numbers.
x = 40
You need 40 pints at 45% and (110-40= 70) 70 pints at 100%